# Large Language Models for Time Series: A Survey

### v1

@@ Abstract
Large Language Models (LLMs) have seen significant use in domains such as natural language processing and computer vision. Going beyond text, image and graphics, LLMs present a significant potential for analysis of time series data, benefiting domains such as climate, IoT, healthcare, traffic, audio and finance. This survey paper provides an in-depth exploration and a detailed taxonomy of the various methodologies employed to harness the power of LLMs for time series analysis. We address the inherent challenge of bridging the gap between LLMs’ original text data training and the numerical nature of time series data, and explore strategies for transferring and distilling knowledge from LLMs to numerical time series analysis. We detail various methodologies, including (1) direct prompting of LLMs, (2) time series quantization, (3) alignment techniques, (4) utilization of the vision modality as a bridging mechanism, and (5) the combination of LLMs with tools. Additionally, this survey offers a comprehensive overview of the existing multimodal time series and text datasets and delves into the challenges and future opportunities of this emerging field. We maintain an up-to-date Github repository [^1] https://github.com/xiyuanzh/awesome-llm-time-series which includes all the papers and datasets discussed in the survey.
@@ 1 Introduction
Figure 2: Left: Taxonomy of LLMs for time series analysis (prompting, quantization, alignment which is further categorized into two groups as detailed in Figure 4, vision as bridge, tool integration). For each category, key distinctions are drawn in comparison to the standard LLM pipeline shown at the top of the figure. Right: We present representative works for each category, sorted by their publication dates. The use of arrows indicates that later works build upon earlier studies. Dark(light)-colored boxes represent billion(million)-parameter models. Icons to the left of the text boxes represent the application domains of domain-specific models, with icons’ meanings illustrated in Figure 1.
@@
In recent years, Large Language Models (LLMs) have gained substantial attention particularly in the fields of Natural Language Processing (NLP) and Computer Vision (CV). Prominent models such as GPT-4 (OpenAI, 2023) have transformed the landscape of text processing by offering unprecedented accuracy in tasks such as text generation, translation, sentiment analysis, question answering and summarization. In the CV domain, LLMs have also facilitated advancements in image recognition, object detection, and generative tasks, leading to more intelligent and capable visual systems (Song et al., 2023). Inspired by these successes, researchers are now exploring the potential of LLMs in the realm of time series analysis, expecting further breakthroughs, as shown in Figure 1. While several surveys offer a broad perspective on large models for time series in general (Jin et al., 2023b; Ma et al., 2023), these do not specifically focus on LLMs or the key challenge of bridging modality gap, which stems from LLMs being originally trained on discrete textual data, in contrast to the continuous numerical nature of time series.
@@
Our survey uniquely contributes to the existing literature by emphasizing how to bridge such modality gap and transfer knowledge from LLMs for time series analysis. Our survey also covers more diverse application domains, ranging from climate, Internet of Things (IoT), to healthcare, traffic management, and finance. Moreover, certain intrinsic properties of time series, like continuity, auto-regressiveness, and dependency on the sampling rate, are also shared by audio, speech, and music data. Therefore, we also present representative LLM-based works from these domains to explore how we can use LLMs for other types of time series. We present a comprehensive taxonomy by categorizing these methodologies into five distinct groups, as shown in Figure 2. If we outline typical LLM-driven NLP pipelines in five stages - input text, tokenization, embedding, LLM, output - then each category of our taxonomy targets one specific stage in this pipeline. Specifically, (i) Prompting (input stage) treats time series data as raw text and directly prompts LLMs with time series; (ii) Time Series Quantization (tokenization stage) discretizes time series as special tokens for LLMs to process; (iii) Alignment (embedding stage) designs time series encoder to align time series embeddings with language space; (iv) Vision as Bridge (LLM stage) connects time series with Vision-Lanuage Models (VLM) by employing visual representations as a bridge; (v) Tool Integration (output stage) adopts language models to output tools to benefit time series analysis. Beyond this taxonomy, our survey also compiles an extensive list of existing multimodal datasets that incorporate both time series and text. We conclude our paper by discussing future research directions in this emerging and promising field.
@@ 3.2 Quantization
Based on VQ-VAE, Auto-TTE (Chung et al., 2023) quantizes ECGs into discrete formats and generates 12-lead ECG signals conditioned on text reports. DeWave (Duan et al., 2023) adapts VQ-VAE to derive discrete codex encoding and aligns it with pre-trained BART for open-vocabulary EEG-to-text translation tasks. TOTEM (Anonymous, 2023c) also quantizes time series through VQ-VAE as input to Transformers for multiple downstream applications such as forecasting, classification, and translation. In the audio domain, UniAudio (Yang et al., 2023) tokenizes different types of target audio using Residual Vector Quantization (RVQ) (Zeghidour et al., 2021) (a hierarchy of multiple vector quantizers) and supports 11 audio generation tasks. VioLA (Wang et al., 2023a) unifies various crossmodal tasks involving speech and text by converting speech utterances to discrete tokens through RVQ. AudioGen (Kreuk et al., 2022) learns discrete audio representations using vector quantization layers and generates audio samples conditioned on text inputs.
@@
Discrete Indices from Other Techniques: Apart from the aforementioned time-domain quantization, FreqTST (Anonymous, 2023b) utilizes frequency spectrum as a common dictionary to discretize time series into frequency units with weights for downstream forecasting task.
@@
3.3 Alignment
(a) Alignment by similarity matching (Type one).
(b) Alignment with large language models as backbones (Type two), where the output could be time series (e.g., forecasting) or text (e.g., EEG-to-text) depending on the downstream tasks.
Figure 4: Two types of alignment based methods.
The third type of works trains a separate encoder for time series, and aligns the encoded time series to the semantic space of language models. These works can be further categorized into two groups based on their specific alignment strategies, as illustrated in Figure 4.
@@ 3.3 Alignment
Table 2: Summary of five major categories of aligning LLMs for time series analysis, including their respective subcategories, representative works, mathematical formulations, advantages and limitations. $q$ and $\mathbf{x}_{v}$ represent text-based quantization process and image data.
@@
Alignment Similarity Match ETP (Liu et al., 2023a) $\mathbf{y}=g_{\phi}(\mathbf{x}_{s})$ align semantics of complicated
@@
For instance, ETP (Liu et al., 2023a; Li et al., 2023a) integrates contrastive learning based pre-training to align electrocardiography (ECG) signals with textual reports. King et al. (2023) use similar contrastive framework to align 17 clinical measurements collected in Intensive Care Unit (ICU) to their corresponding clinical notes. TEST (Sun et al., 2023) uses contrastive learning to generate instance-wise, feature-wise, and text-prototype-aligned time series embeddings to align with text embeddings. TENT (Zhou et al., 2023b) aligns text embeddings with IoT sensor signals through a unified semantic feature space using contrastive learning. JoLT (Cai et al., 2023) utilizes Querying Transformer (Q-Former) (Li et al., 2023b) optimized with contrastive loss to align the time series and text representations.
@@
Similarity Matching through Other Losses: Apart from contrastive loss, other loss functions are also employed to optimize similarity matching between time series embeddings and text embeddings. ECG-LLM (Qiu et al., 2023) aligns the distribution between ECG and language embedding from ECG statements with an Optimal Transport based loss function to train an ECG report generation model. MTAM (Han et al., 2022) uses various alignment techniques, such as Canonical Correlation Analysis and Wasserstein Distance, as loss functions to align electroencephalography (EEG) features with their corresponding language descriptions.
@@
LLMs as Backbones: The second type of alignment method directly uses large language models as backbones following time series embedding layers. EEG-to-Text (Wang and Ji, 2022) feeds EEG embeddings to pre-trained BART for open vocabulary EEG-To-Text decoding and EEG-based sentiment classification. GPT4TS (Zhou et al., 2023a) uses patching embeddings (Nie et al., 2022) as input to frozen pre-trained GPT-2 where the positional embedding layers and self-attention blocks are retained during time series fine-tuning. The method provides a unified framework for seven time series tasks, including few-shot or zero-shot learning. Following GPT4TS, researchers further incorporated seasonal-trend decomposition (TEMPO (Cao et al., 2023)), two-stage fine-tuning (LLM4TS (Chang et al., 2023)), domain descriptions (UniTime (Liu et al., 2023e)), graph attention mechanism (GATGPT (Chen et al., 2023)), and spatial-temporal embedding module (ST-LLM (Liu et al., 2024)). Time-LLM (Jin et al., 2023a) reprograms time series data into text prototypes as input to LLaMA-7B. It also provides natural language prompts such as domain expert knowledge and task instructions to augment input context. Lag-Llama (Rasul et al., 2023) builds univariate probabilistic time series forecasting model based on LLaMA architecture. In the audio, speech and music domains, researchers have also designed dedicated encoders to embed speech (WavPrompt  (Gao et al., 2022), Speech LLaMA (Lakomkin et al., 2023))), music (MU-LLaMA (Liu et al., 2023c)), and general audio inputs (LTU (Gong et al., 2023), SALMONN (Tang et al., 2023)), and feed the embeddings to large language models.
@@ 3.4 Vision as Bridge
Paired Data: ImageBind (Girdhar et al., 2023) uses image-paired data to bind six modalities (images, text, audio, depth, thermal, and Inertial Measurement Unit (IMU) time series) and learn a joint embedding space, enabling new emergent alignments and capabilities. PandaGPT (Su et al., 2023) further combines the multimodal encoders from ImageBind and LLMs to enable visual and auditory instruction-following capabilities. IMU2CLIP (Moon et al., 2022) aligns IMU time series with video and text, by projecting them into the joint representation space of Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021). AnyMAL (Moon et al., 2023) builds upon IMU2CLIP by training a lightweight adapter to project the IMU embeddings into the text token embedding space of LLaMA-2-70B. It is also capable of transforming data from other modalities, such as images, videos, audio, into the same text embedding space.
@@ 3.5 Tool
MoAT [^5] https://openreview.net/pdf?id=uRXxnoqDHH (Anonymous, 2023a) $6$ datasets, $2$K timesteps in total text, time series forecasting
@@ 4 Comparison within the Taxonomy
Data: When no training data is available and the objective is to apply LLM for time series in an zero-shot fashion, it is preferable to use prompting-based methods. This is because direct prompting enables the utilization of pre-trained language models’ inherent capabilities without fine-tuning. However, representing numbers as strings can diminish the semantic value intrinsically tied to numerical data. Therefore, with adequate training data, quantization or alignment-based methods become more advantageous. As shown in Figure 2, these two categories are the most extensively studied ones in existing literature. Furthermore, if time series data can be interpreted or associated with visual representations, these representations can be incorporated to utilize the intrinsic knowledge embedded in the vision modality or pre-trained vision-language models.
Model: Prompting and tool integration methods tend to apply billion-parameter models as they often apply off-the-self LLMs without architectural modifications. By contrast, alignment and quantization methods vary from million to billion-parameter models, depending on the specific application requirements and available computational resources.
Efficiency: Prompting-based methods are not efficient for numerical data with high precision, as well as multivariate time series as it requires transforming each dimension into separate univariate time series, resulting in extremely long input. They are also less efficient for long-term predictions due to the computational demands of generating long sequences. These methods are more effective when dealing with simple numerical data that is richly interwoven with textual information, such as opening and closing stock prices in financial news articles. By contrast, quantization and alignment methods are more efficient to handle long sequences, as time series are typically down-sampled or segmented into patches before feeding into large language models.
Optimization: Depending on the specific discretization technique, quantization based method may require a two-stage training process (such as first training the VQ-VAE model), which may result in sub-optimal performance compared with that achieved through end-to-end training in alignment methods. Using large language models as indirect tools empowers LLMs with more capabilities to manage numerical data, but also raises the level of complexity to optimize both LLMs and other components in an end-to-end fashion. Therefore, existing works of tool integration typically employ off-the-shelf LLMs without further fine-tuning.
@@ 5 Multimodal Datasets
Applying LLMs for time series benefits from the availability of multimodal time series and text data. In this section, we introduce representative multimodal time series and text datasets organized by their respective domains (Table 3).
@@
Finance: PIXIU (Xie et al., 2023b) presents multi-task and multi-modal instruction tuning data in the financial domain with 136K data samples. It contains both financial natural language understanding and prediction tasks, and covers 9 datasets of multiple modalities such as text and time series. MoAT (Anonymous, 2023a) constructs multimodal datasets with textual information paired with time series for each timestep, such as news articles extracted with relevant keywords, mostly covering finance related domains such as fuel, metal, stock and bitcoin.

### v3

@@ Abstract
Large Language Models (LLMs) have seen significant use in domains such as natural language processing and computer vision. Going beyond text, image and graphics, LLMs present a significant potential for analysis of time series data, benefiting domains such as climate, IoT, healthcare, traffic, audio and finance. This survey paper provides an in-depth exploration and a detailed taxonomy of the various methodologies employed to harness the power of LLMs for time series analysis. We address the inherent challenge of bridging the gap between LLMs’ original text data training and the numerical nature of time series data, and explore strategies for transferring and distilling knowledge from LLMs to numerical time series analysis. We detail various methodologies, including (1) direct prompting of LLMs, (2) time series quantization, (3) aligning techniques, (4) utilization of the vision modality as a bridging mechanism, and (5) the combination of LLMs with tools. Additionally, this survey offers a comprehensive overview of the existing multimodal time series and text datasets and delves into the challenges and future opportunities of this emerging field. We maintain an up-to-date Github repository [^1] https://github.com/xiyuanzh/awesome-llm-time-series which includes all the papers and datasets discussed in the survey.
@@ 1 Introduction
Figure 2: Left: Taxonomy of LLMs for time series analysis (prompting, quantization, aligning which is further categorized into two groups as detailed in Figure 4, vision as bridge, tool integration). For each category, key distinctions are drawn in comparison to the standard LLM pipeline shown at the top of the figure. Right: We present representative works for each category, sorted by their publication dates. The use of arrows indicates that later works build upon earlier studies. Dark(light)-colored boxes represent billion(million)-parameter models. Icons to the left of the text boxes represent the application domains of domain-specific models, with icons’ meanings illustrated in Figure 1.
@@
In recent years, Large Language Models (LLMs) have gained substantial attention particularly in the fields of Natural Language Processing (NLP) and Computer Vision (CV). Prominent models such as GPT-4 (OpenAI, 2023) have transformed the landscape of text processing by offering unprecedented accuracy in tasks such as text generation, translation, sentiment analysis, question answering and summarization. In the CV domain, Large Multimodal Models (LMMs) have also facilitated advancements in image recognition, object detection, and generative tasks, leading to more intelligent and capable visual systems (Song et al., 2023). Inspired by these successes, researchers are now exploring the potential of LLMs in the realm of time series analysis, expecting further breakthroughs, as shown in Figure 1. While several surveys offer a broad perspective on large models for time series in general (Jin et al., 2023b; Ma et al., 2023), these do not specifically focus on LLMs or the key challenge of bridging modality gap, which stems from LLMs being originally trained on discrete textual data, in contrast to the continuous numerical nature of time series.
@@
Our survey uniquely contributes to the existing literature by emphasizing how to bridge such modality gap and transfer knowledge from LLMs for time series analysis. Our survey also covers more diverse application domains, ranging from climate, Internet of Things (IoT), to healthcare, traffic management, and finance. Moreover, certain intrinsic properties of time series, like continuity, auto-regressiveness, and dependency on the sampling rate, are also shared by audio, speech, and music data. Therefore, we also present representative LLM-based works from these domains to explore how we can use LLMs for other types of time series. We present a comprehensive taxonomy by categorizing these methodologies into five distinct groups, as shown in Figure 2. If we outline typical LLM-driven NLP pipelines in five stages - input text, tokenization, embedding, LLM, output - then each category of our taxonomy targets one specific stage in this pipeline. Specifically, (i) Prompting (input stage) treats time series data as raw text and directly prompts LLMs with time series; (ii) Time Series Quantization (tokenization stage) discretizes time series as special tokens for LLMs to process; (iii) Aligning (embedding stage) designs time series encoder to align time series embeddings with language space; (iv) Vision as Bridge (LLM stage) connects time series with Vision-Lanuage Models (VLM) by employing visual representations as a bridge; (v) Tool Integration (output stage) adopts language models to output tools to benefit time series analysis. Beyond this taxonomy, our survey also compiles an extensive list of existing multimodal datasets that incorporate both time series and text. We conclude our paper by discussing future research directions in this emerging and promising field.
@@ 3.2 Quantization
Based on VQ-VAE, Auto-TTE (Chung et al., 2023) quantizes ECGs into discrete formats and generates 12-lead ECG signals conditioned on text reports. DeWave (Duan et al., 2023) adapts VQ-VAE to derive discrete codex encoding and aligns it with pre-trained BART for open-vocabulary EEG-to-text translation tasks. TOTEM (Talukder and Gkioxari, 2023) also quantizes time series through VQ-VAE as input to Transformers for multiple downstream applications such as forecasting, classification, and translation. In the audio domain, UniAudio (Yang et al., 2023) tokenizes different types of target audio using Residual Vector Quantization (RVQ) (Zeghidour et al., 2021) (a hierarchy of multiple vector quantizers) and supports 11 audio generation tasks. VioLA (Wang et al., 2023a) unifies various crossmodal tasks involving speech and text by converting speech utterances to discrete tokens through RVQ. AudioGen (Kreuk et al., 2022) learns discrete audio representations using vector quantization layers and generates audio samples conditioned on text inputs.
@@
Discrete Indices from Other Techniques: Apart from the aforementioned time-domain quantization, FreqTST (Li et al., 2023b) utilizes frequency spectrum as a common dictionary to discretize time series into frequency units with weights for downstream forecasting task. Chronos (Ansari et al., 2024) quantizes real-valued time series into discrete bins, and optimizes existing language model architectures on these tokenized time series via the cross-entropy loss.
@@
3.3 Aligning
(a) Aligning by similarity matching (Type one).
(b) Aligning with large language models as backbones (Type two), where the output could be time series (e.g., forecasting) or text (e.g., EEG-to-text) depending on the downstream tasks.
Figure 4: Two types of aligning based methods.
The third type of works trains a separate encoder for time series, and aligns the encoded time series to the semantic space of language models. These works can be further categorized into two groups based on their specific aligning strategies, as illustrated in Figure 4.
@@ 3.3 Aligning
Table 2: Summary of five major categories of applying LLMs for time series analysis, including their respective subcategories, representative works, mathematical formulations, advantages and limitations. $q$ and $\mathbf{x}_{v}$ represent text-based quantization process and image data.
@@
Aligning Similarity Match ETP (Liu et al., 2023a) $\mathbf{y}=g_{\phi}(\mathbf{x}_{s})$ align semantics of complicated
@@
For instance, ETP (Liu et al., 2023a; Li et al., 2023a) integrates contrastive learning based pre-training to align electrocardiography (ECG) signals with textual reports. King et al. (2023) use similar contrastive framework to align 17 clinical measurements collected in Intensive Care Unit (ICU) to their corresponding clinical notes. TEST (Sun et al., 2023) uses contrastive learning to generate instance-wise, feature-wise, and text-prototype-aligned time series embeddings to align with text embeddings. TENT (Zhou et al., 2023b) aligns text embeddings with IoT sensor signals through a unified semantic feature space using contrastive learning. JoLT (Cai et al., 2023) utilizes Querying Transformer (Q-Former) (Li et al., 2023c) optimized with contrastive loss to align the time series and text representations.
@@
Similarity Matching through Other Losses: Apart from contrastive loss, other loss functions are also employed to optimize similarity matching between time series embeddings and text embeddings. ECG-LLM (Qiu et al., 2023) aligns the distribution between ECG and language embedding from ECG statements with an Optimal Transport based loss function to train an ECG report generation model. MTAM (Han et al., 2022) uses various aligning techniques, such as Canonical Correlation Analysis and Wasserstein Distance, as loss functions to align electroencephalography (EEG) features with their corresponding language descriptions.
@@
LLMs as Backbones: The second type of aligning method directly uses large language models as backbones following time series embedding layers. EEG-to-Text (Wang and Ji, 2022) feeds EEG embeddings to pre-trained BART for open vocabulary EEG-To-Text decoding and EEG-based sentiment classification. GPT4TS (Zhou et al., 2023a) uses patching embeddings (Nie et al., 2022) as input to frozen pre-trained GPT-2 where the positional embedding layers and self-attention blocks are retained during time series fine-tuning. The method provides a unified framework for seven time series tasks, including few-shot or zero-shot learning. Following GPT4TS, researchers further incorporated seasonal-trend decomposition (TEMPO (Cao et al., 2023)), two-stage fine-tuning (LLM4TS (Chang et al., 2023)), domain descriptions (UniTime (Liu et al., 2023e)), graph attention mechanism (GATGPT (Chen et al., 2023)), and spatial-temporal embedding module (ST-LLM (Liu et al., 2024)). Time-LLM (Jin et al., 2023a) reprograms time series data into text prototypes as input to LLaMA-7B. It also provides natural language prompts such as domain expert knowledge and task instructions to augment input context. Lag-Llama (Rasul et al., 2023) builds univariate probabilistic time series forecasting model based on LLaMA architecture. In the audio, speech and music domains, researchers have also designed dedicated encoders to embed speech (WavPrompt  (Gao et al., 2022), Speech LLaMA (Lakomkin et al., 2023))), music (MU-LLaMA (Liu et al., 2023c)), and general audio inputs (LTU (Gong et al., 2023), SALMONN (Tang et al., 2023)), and feed the embeddings to large language models.
@@ 3.4 Vision as Bridge
Paired Data: ImageBind (Girdhar et al., 2023) uses image-paired data to bind six modalities (images, text, audio, depth, thermal, and Inertial Measurement Unit (IMU) time series) and learn a joint embedding space, enabling new emergent capabilities. PandaGPT (Su et al., 2023) further combines the multimodal encoders from ImageBind and LLMs to enable visual and auditory instruction-following capabilities. IMU2CLIP (Moon et al., 2022) aligns IMU time series with video and text, by projecting them into the joint representation space of Contrastive Language-Image Pre-training (CLIP) (Radford et al., 2021). AnyMAL (Moon et al., 2023) builds upon IMU2CLIP by training a lightweight adapter to project the IMU embeddings into the text token embedding space of LLaMA-2-70B. It is also capable of transforming data from other modalities, such as images, videos, audio, into the same text embedding space.
@@ 3.5 Tool
MoAT [^5] https://openreview.net/pdf?id=uRXxnoqDHH (Lee et al., 2023) $6$ datasets, $2$K timesteps in total text, time series forecasting
@@ 4 Comparison within the Taxonomy
Data: When no training data is available and the objective is to apply LLM for time series in an zero-shot fashion, it is preferable to use prompting-based methods. This is because direct prompting enables the utilization of pre-trained language models’ inherent capabilities without fine-tuning. However, representing numbers as strings can diminish the semantic value intrinsically tied to numerical data. Therefore, with adequate training data, quantization or aligning-based methods become more advantageous. As shown in Figure 2, these two categories are the most extensively studied ones in existing literature. Furthermore, if time series data can be interpreted or associated with visual representations, these representations can be incorporated to utilize the intrinsic knowledge embedded in the vision modality or pre-trained vision-language models.
Model: Prompting and tool integration methods tend to apply billion-parameter models as they often apply off-the-self LLMs without architectural modifications. By contrast, aligning and quantization methods vary from million to billion-parameter models, depending on the specific application requirements and available computational resources.
Efficiency: Prompting-based methods are not efficient for numerical data with high precision, as well as multivariate time series as it requires transforming each dimension into separate univariate time series, resulting in extremely long input. They are also less efficient for long-term predictions due to the computational demands of generating long sequences. These methods are more effective when dealing with simple numerical data that is richly interwoven with textual information, such as opening and closing stock prices in financial news articles. By contrast, quantization and aligning methods are more efficient to handle long sequences, as time series are typically down-sampled or segmented into patches before feeding into large language models.
Optimization: Depending on the specific discretization technique, quantization based method may require a two-stage training process (such as first training the VQ-VAE model), which may result in sub-optimal performance compared with that achieved through end-to-end training in aligning methods. Using large language models as indirect tools empowers LLMs with more capabilities to manage numerical data, but also raises the level of complexity to optimize both LLMs and other components in an end-to-end fashion. Therefore, existing works of tool integration typically employ off-the-shelf LLMs without further fine-tuning.
@@ 5 Multimodal Datasets
Applying LLMs for time series benefits from the availability of multimodal time series and text data. In this section, we introduce representative multimodal time series and text datasets organized by their respective domains (Table 3). We list additional multimodal datasets in our Github repository [^12] https://github.com/xiyuanzh/awesome-llm-time-series.
@@
Finance: PIXIU (Xie et al., 2023b) presents multi-task and multi-modal instruction tuning data in the financial domain with 136K data samples. It contains both financial natural language understanding and prediction tasks, and covers 9 datasets of multiple modalities such as text and time series. MoAT (Lee et al., 2023) constructs multimodal datasets with textual information paired with time series for each timestep, such as news articles extracted with relevant keywords, mostly covering finance related domains such as fuel, metal, stock and bitcoin.
@@
These datasets offer valuable benchmarks for multimodal time series and text analysis. These contain both time series focused tasks, including classification, which is evaluated using accuracy and macro-F1 scores, and forecasting, which utilizes metrics such as MSE, MAE, RMSE, and MAPE, as well as NLP focused tasks such as captioning, question answering, and translation, assessed through BLEU, ROUGE, METEOR, and EM scores, among others.